OpenStudy (anonymous):

Which equation is related to this equation? k + 26 = 61 A. k = 26 – 61 B. k = 61 – 26 C. k = 61 ÷ 26 D. k = 61 • 26

OpenStudy (anonymous):

@thelaw2 A or B

OpenStudy (anonymous):

Im pretty sure A

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok so k + 26 = 61 A. k = 26 – 61 B. k = 61 – 26 What do you need to do first? You get the k alone to find out what k is correct? So you minus 26 on both sides to make it even. So now it's k + 0 = 61 - 26.

OpenStudy (anonymous):

Why is A wrong? Because you're subtracting 61 from 26 instead of subtracting 26 from 61.

OpenStudy (anonymous):

So B That makes sense

OpenStudy (anonymous):

Yup